OpenStudy (anonymous):

What is the simplified solution to five over eighteen divided by 5 over twelve? one over three two over three seventeen over twenty-three twenty-five over two hundred sixteen

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

\[5/18 \div 5/12\] =5/18 x 12/5 =12/18 =2/3